OpenStudy (anonymous):

Simplify: 3/x + 2/x+2 / 3/x+2 - 2x

OpenStudy (anonymous):

@johnweldon1993 it's two fractions on top 3 over x plus 2 over x then divided by 2 fractions on the bottom 3 over x +2 minus 2x

OpenStudy (johnweldon1993):

huh? lol \[\huge \frac{\frac{3}{x} + \frac{2}{x+2}}{\frac{3}{x + 2} - 2x}\] lol like that?

OpenStudy (anonymous):

Everything is correct except for 2x is 2/x

OpenStudy (johnweldon1993):

\[\huge \frac{\frac{3}{x} + \frac{2}{x+2}}{\frac{3}{x + 2} - \frac{2}{x}}\]

OpenStudy (anonymous):

Yep that's it

OpenStudy (anonymous):

:D

OpenStudy (johnweldon1993):

This one is annoying I cant quite break it down easily

OpenStudy (johnweldon1993):

What are your answer choices for this one? lol

OpenStudy (vivek3461):

\[\frac{ 5x + 6 }{ x-4 }\]

OpenStudy (anonymous):

Numerator: \[ \frac{3}{x} + \frac{2}{x + 2} = \frac{3(x + 2) + 2(x)}{x(x + 2)} = \frac{5x + 6}{x(x + 2)}\] Denominator: \[ \frac{3}{x + 2} - \frac{2}{x} = \frac{3(x) - 2(x + 2)}{(x + 2)x} = \frac{x - 4}{x(x + 2)} \] Now we'll find \(\frac{\text{numerator}}{\text{denominator}}\). Notice that the subdenominators x(x + 2) are the same. They cancel. Left we've got \[ \frac{5x + 6}{x - 4} \]